# Statistical testing of random number generators and their improvement using randomness extraction

### v1

@@ V.6.1 Results
  • Randomness that has a small amount of min-entropy only can pass statistical tests successfully. This is somewhat unsurprising, since many cryptographically secure PRNGs exist, but we find it interesting to comment on nonetheless. The total entropy of the final output of the processed LFSR output is upper bounded by $10007+32$ (the seed length of the extractor plus the seed length of the 32-bit LFSR), in the 50Gbit of processed output generated, i.e. a true min-entropy rate of, at most, $\alpha=(10007+32)/(5\times 10^{10})<10^{-5}$.
Our results at this level are disappointing, in the sense that the successful test results mean that we will not be able to distinguish the next levels (3 and 4) from level 2 from a statistical perspective – for example that level 3 is strictly better than level 2. It would be interesting to find non-trivial examples where the output of a seeded extractor fails statistical tests because of a seed generated in a way that is not independent or near-uniform. Unfortunately, we could only find artificial examples (i.e. when all seed bits are the same) that get detected by our statistical testing.
V.7 Level 3: Two-source extraction
Seeded extraction (level 2) requires an independent string of (near-)perfect randomness as an initial resource, which is hard to justify and leads to a circularity: one needs near-perfect randomness to generate more of it. Two-source extraction relaxes this requirement, allowing the second string to be only weakly random, in the sense that it has some min-entropy and/or only a relaxed notion of independence [^11] For example, the case of using a two-source extractor secure in the Markov model arnon2015quantum , where the two input sources can be correlated through a common cause, or if the sources may have bounded coordination, cross-influence or mutual information ball2022randomness . – although in this work we calculate our two-source extractor parameters based on standard independence between the two input sources. Two-source extractors can be used as seeded extractors, simply by assuming that one of the input strings is already near-perfect and independent, therefore level 3 is strictly a relaxation of the assumptions of level 2.
Definition 10 (Two-source randomness extractor).
A two-source randomness extractor is a function $\mathsf{Ext_{2}}:\{0,1\}^{n_{1}}\times\{0,1\}^{n_{2}}\rightarrow\{0,1\}^{m}$ such that, for statistically independent random variables $X\in\{0,1\}^{n_{1}}$ and $Y\in\{0,1\}^{n_{2}}$ with min-entropy’s $\mathrm{H}_{\infty}(X)\geq k_{1}$ and $\mathrm{H}_{\infty}(Y)\geq k_{2}$ respectively,
$\Delta(\mathsf{Ext_{2}}(X,Y),\mathrm{U}_{m})\leq\epsilon$ (13)
where $\mathrm{U}_{m}$ is the uniform variable on $\{0,1\}^{m}$.
In other words, a two-source extractor is a weakly randomised function that maps a random variable $X$ to a new variable $\mathsf{Ext_{2}}(X,Y)$ that is near-perfect.
Definition 11 (Strong two-source extractor).
A two-source randomness extractor is said to be strong in the input $Y$ if the function $\mathsf{Ext_{2}}$ is such that
$\Delta([\mathsf{Ext_{2}}(X,Y),Y],[\mathrm{U}_{m}Y])\leq\epsilon$ (14)
where [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] denotes the concatenation of random variables and $\mathrm{U}_{m}$ is the uniform variable on $\{0,1\}^{m}$.
Strong two-source extractors, like strong seeded extractors, allow for one input source to be used in multiple extraction rounds.
Figure 7: The set-up for two-source extraction. In this case, the initial output of the RNG only needs to have some min-entropy and extraction requires an additional bit string which is weakly random only, in the sense that it also has min-entropy.
From the Cryptomite library, we again use the Circulant extractor cryptomite , but this time, as a strong two-source extractor. Other extractors in Cryptomite can be used too, but, since the Circulant extractor offers the best parameters and efficiency, we use it in our implementation. For full details, we refer the reader to cryptomite . Two-source extraction requires a second input source with min-entropy above some threshold based on the specific two-source extractor construction. For the Circulant extractor, this requirement is that the sum of the min-entropy rates of the two weak inputs is at least 1. $X_{i}$ is the initial RNG output blocks and $Y$ is the additional weakly random input (which we sometimes call the weak seed) and, as in level 2, we set $|Y|=|X_{i}|+1=10007$.
To generate $Y$, we again use the NIST Randomness Beacon, but, in this case we minimise the amount of entropy we assume it contains instead of assuming it has full entropy as in level 2. This change in the assumption increases the likelihood that the assumption holds in practice. The output length of the Circulant extractor is roughly $(\alpha_{\mathsf{NIST}}+\alpha_{\mathsf{RNG}}-1)|Y|$, which we impose by adjusting the min-entropy rate assumption of the NIST Randomness Beacon as $\alpha_{\mathsf{NIST}}$, as
$\alpha_{\mathsf{NIST}}=1.02-\alpha_{\mathsf{RNG}},$ (15)
where $\alpha_{\mathsf{RNG}}$ is the min-entropy rate of the initial RNG (found in Section IV). We use 1.02 instead of 1 to account spurious terms in the parameter calculation that reduce the output length, see cryptomite for the explicit calculation of these penalty terms. In other words, we use the computed min-entropy rate of the RNG under study to minimise the assumption about the second source’s min-entropy rate, whilst imposing a non-trivial output length from the extractor.
The processed output is then generated in two steps. (1) using the Circulant extractor as a two-source extractor on the two input strings $X_{1}$ and $Y$, we generate a (near-)perfect output which will be the seed in the next step. (2) use this seed in multiple Circulant seeded extractions on $X_{i\geq 2}$. The multiple outputs of the seeded extractor are concatenated together to obtain a final output of $5\times 10^{10}$ bits. In other words, the concatenation of the two-source and seeded extractors together form a two-source extractor with advantageous parameters. Therefore, the final output for statistical testing is given by:
$\mathsf{Output}=\left[\mathsf{Ext_{s}^{Circulant}}(X_{2},S),\mathsf{Ext_{s}^{Circulant}}(X_{3},S),\ldots,\mathsf{Ext_{s}^{Circulant}}(X_{n},S)\right],\quad\quad S=\mathsf{Ext_{2}^{Circulant}}(X_{1},Y),$ (16)
where [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] denotes the concatenation of random variables and the extractor round with input $X_{i}$ has error $\epsilon_{\mathsf{ext_{i}}}$. The total error of the final output is $\epsilon_{\mathsf{total}}=\epsilon_{\mathsf{est}}+\epsilon_{\mathsf{ext_{1}}}+\sum_{j=2}^{n}\epsilon_{\mathsf{ext_{j}}}\leq 2^{-32}$. A proof that a strong two-source extractor and strong seeded extractor can be composed into $\mathsf{Ext_{s}^{Circulant}}(X_{i>1},S)$, for $S$ the output of a two-source extractor (right hand side of Equation 16) can be found in vadhan2012pseudorandomness Section 6.3. This, combined with the fact that composed output error is the sum of each of the individual extractor errors (in cryptomite ) allows us to calculate $\epsilon_{\mathsf{total}}$.
V.7.1 Results
RNG
NIST
Diehard
ENT
SmallCrush
Alphabit
Rabbit
PractRand
32-bit LFSR 0 0 (6) 0 0 2 1 0 (8)
RDSEED 0 0 (4) 0 0 0 3 0 (5)
IDQ Quantis 0 0 (3) 0 0 0 1 0 (5)
@@ V.7.1 Results
Table 7: This table gives the sum of tests failed for $5\times 10$Gbit samples from each RNG, after strong two-source extraction taking the RNG as one weak source and randomness from the NIST Randomness Beacon as the second. In cells with multiple entries, failed tests are on the left and suspicious tests (when applicable) are on the right in parenthesis. Full results can be found in Section E.2.
@@
Our results show that all RNGs extracted at level 3 are successful from a statistical perspective, like in the seeded extraction case (level 2). In the appendices, we implement a variant of level 3 (two-source extraction) where all input strings are drawn from the initial RNG and there is no randomness from an alternative RNG, i.e. rewriting the $\mathsf{Output}$ in Equation 16 using $Y=X_{0}$, where $X_{0}$ is another output block from the initial RNG. In this regime, for near-perfect randomness to be generated, each block produced by the initial RNG must be independent from one another (as well as have block min-entropy). Even in this case, the results were successful statistically. Full explanation and results can be found in Section E.1.
V.8 Level 4: Physical randomness extraction
@@
Two-source extraction (level 3) allows for the generation of near-perfect randomness if two, weakly random but independent strings of randomness are available. In the final level, we consider post-processing with a physical randomness extractor. This level is called physical because it requires a quantum device, in addition to the initial RNG, while the other levels only required mathematical algorithms to perform extraction. At a high level, the role of this additional hardware is to provide a second string of random numbers, whilst making minimal assumptions only.
@@
Adding quantum hardware may initially seem to imply introducing numerous assumptions, however, following the device-independent approach, this hardware can in principle be treated as an untrusted black box (which could even have been built by an adversary, so long as it can be shielded once in use and meets some minimal requirements). We call the added assumptions minimal because they are either fundamental to physics – e.g. information cannot travel faster than light speed – or no cryptography can ever be done without them – e.g. the devices are shielded (there are no backdoors). This is made possible by the development of device-independent protocols, which rely on Bell tests brunner2014bell . The idea is to use the initial RNG to generate random challenges for the quantum device, and then studying its response. With ideal (noiseless) devices, this approach can be used to self-test the inner functioning of the device, i.e. one can uniquely identify the implemented quantum states and measurements from the observed challenge-response statistics alone. For real (noisy) devices, this approach can be used to bound the adversary’s guessing power, and thus guarantee min-entropy, over the device’s outputs or responses. For a review on the subject, together with its minimal assumptions (called loopholes), we refer the reader to acin2016certified . This approach crucially relies on quantum resources, which have this self-testing property, and has no classical analogue. See Figure 8 for an illustration.
Figure 8: The set up for level 4: physical randomness extraction. The initial RNG is used twice: first to generate challenges to the quantum device, and second, to provide an extra bit string as input to a two-source extractor. The role of the quantum device is to provide an additional source of randomness. The device-independent protocol is performed by using the challenge-response behaviour of the device to obtain a lower bound on the amount of randomness in the device’s responses (without characterising the device itself). The second bit string of the initial RNG and the responses from the quantum device form the two input strings to a two-source extractor, implemented as in level 3.
@@
Today, quantum devices that are capable of running device-independent protocols are extremely hard to build (they require the ability to perform a loophole free Bell Test brunner2014bell ) and exist as experiments on lab benches only. Because of this, more practical implementations have been developed in which a few, well justified, assumptions are added (to the minimal ones). The resulting protocol thus still has comparatively fewer assumptions than adding standard hardware, but is not minimal either. Such a semi-device-independent protocol is the one we implement for our physical extraction method at level 4, based on the randomness amplification protocol described in foreman2020practical and implemented on quantum computers. For clarity, the assumptions are:
@@
  • The quantum device is assumed to perform a faithful Bell test. This assumption is well motivated when using particular types of device, such as the quantum computers based on ion-traps that we use – see the discussion in foreman2020practical (Section 6.2, Validity of quantum computers for Bell experiments and added assumptions).
@@
We used the $\mathsf{H}$1-1 Quantinuum ion-trap quantum computer h1-1 as our device to obtain, from its output, a weakly random bit string size of $3.6\times 10^{6}$ bits [^12] This means, because of using the Circulant extractor, the input length of the initial RNG block to the two-source extraction is also $3.6\times 10^{6}$ bits. with min-entropy rate $\alpha_{\mathsf{Q}}\geq 0.518$, certified in the semi-device-independent manner described above. The Circulant extractor requires $\alpha_{\mathsf{Q}}+\alpha_{\mathsf{RNG}}>1$ to give non-vanishing output, implying that the rate of an initial RNG must satisfy $\alpha_{\mathsf{RNG}}>0.482$ to allow for physical extraction with our implementation [^13] This minimum requirement is particularly interesting, since, even if one has access to two identical (and independent) copies of an initial RNG with $\alpha_{\mathsf{RNG}}=0.482+\delta$ for $\delta\in(0,0.18)$, one would be unable to extract from the two (step 3) with today’s implemented extractors. Note that this is not a fundamental limitation, as other two-source extractors allow for one of the strings to have logarithmic min-entropy rate only. However, up to our knowledge no such extractor has been implemented, let alone efficiently, since $0.482+0.482<1$, making the results of this section more interesting. The advantage of using a quantum device, and therefore level 4, is two-fold: a) one gets a rigorous, semi-device-independent, lower bound on a second bit string’s min-entropy and, b) the min-entropy rate of the quantum device is above $0.5$, allowing the extraction from a weak initial output with rate $0.5$ using the Circulant extractor. Note that the min-entropy of the LSFR was too low to perform physical extraction (its min-entropy rate is below $0.482$, see Table 4).
@@
The processed output is then generated in two steps. (1) Generate a (near-)perfect seed using the Circulant extractor as a two-source extractor on the two input strings $X_{1}$, from the initial RNG, and $Y$, from the $\mathsf{H}$1-1 Quantinuum quantum computer. (2) Use this seed in multiple Circulant seeded extractions on $X_{i\geq 2}$ which are concatenated together to obtain a final output of $5\times 10^{10}$ bits. In other words, the concatenation of the two-source and seeded extractors together again form a two-source extractor with advantageous parameters. Therefore, the final output for statistical testing is given by:
$\mathsf{Output}=\left[\mathsf{Ext_{s}^{Circulant}}(X_{2},S),\mathsf{Ext_{s}^{Circulant}}(X_{3},S),\ldots,\mathsf{Ext_{s}^{Circulant}}(X_{n},S)\right],\quad\quad S=\mathsf{Ext_{2}^{Circulant}}(X_{1},Y),$ (17)
@@
where [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] denotes concatenation and the extractor round with input $X_{i}$ has error $\epsilon_{\mathsf{ext_{i}}}$. The total error of the final output is $\epsilon_{\mathsf{total}}=\epsilon_{\mathsf{est}}+\epsilon_{\mathsf{ext_{1}}}+\sum_{j=2}^{n}\epsilon_{\mathsf{ext_{j}}}\leq 2^{-32}$. This last step is similar to that of level 3, where the NIST Randomness Beacon is replaced by the $\mathsf{H}$1-1 Quantinuum quantum computer. The statistical test results are given in the following table.
V.8.1 Results
RNG
NIST
Diehard
ENT
SmallCrush
Alphabit
Rabbit
PractRand
32-bit LFSR - - - - - - -
RDSEED 0 0 (2) 0 0 0 1 0 (3)
IDQ Quantis 0 0 (3) 1 0 0 2 0 (7)
@@ V.8.1 Results
Table 8: This table gives the sum of tests failed for $5\times 10$Gbit samples from level 4. Note: The 32-bit LFSR does not generate any output in this setting, since it’s min-entropy is too low for extraction. In cells with multiple entries, failed tests are on the left and suspicious tests (when applicable) are on the right in parenthesis. Full results can be found in Appendix F.
@@
The statistical test results show, as for level 2 and 3, that the post-processed RNGs perform well at level 4.
V.9 Environmental impact
A consumer-grade laptop consumes roughly 200 W/hour under heavy usage. The total runtime of our test suite with parameter ‘all’ on is 128 minutes, so we estimate the energy used in conducting a single run of statistical testing to be around 0.425kWh (1’530’ [^0] J), which equates to roughly 0.184kg of CO2 (using https://www.epa.gov/energy/greenhouse-gas-equivalencies-calculator). For the entirety of this work, we performed full testing (i.e. using the all setting) 105 times giving a total energy usage of around 44.625kWh (160’650’ [^0] J), and therefore, roughly generated 19.32kg of CO2.
VI Conclusion and Future Work

### v2

@@ 5.2.2 Level 2: Seeded Extraction
  • Randomness that has a small amount of min-entropy only can pass statistical tests successfully. This is somewhat unsurprising as cryptographically secure PRNGs exist, but we find it interesting to comment on nonetheless. The total entropy of the final output of the processed LFSR output is upper-bounded by 10,007 $+32$ (the seed length of the extractor plus the seed length of the 32-bit LFSR), in the 50 Gbit of processed output generated, i.e., a true min-entropy rate of, at most, $\alpha=(10,007+32)/(5\times 10^{10})<10^{-5}$.
@@
Table 6: This table gives the sum of statistical tests failed for $5\times 10$ Gbit samples from each RNG, after a strong seeded extractor has been applied to its initial output. The seed is generated using the NIST Randomness Beacon. In cells with multiple entries, failed tests are on the left and suspicious tests (when applicable) are on the right. Full results can be found in Appendix D.
{adjustwidth}
[…]
Our results at this level are disappointing, in the sense that the successful test results mean that we will not be able to distinguish the next levels (3 and 4) from level 2 from a statistical perspective—for example, that level 3 is strictly better than level 2. It would be interesting to find non-trivial examples where the output of a seeded extractor fails statistical tests because of a seed generated in a way that is not independent or near-uniform. Unfortunately, we could only find artificial examples (i.e., when all seed bits are the same) that were detected by our statistical testing.
5.2.3 Level 3: Two-Source Extraction
Seeded extraction (level 2) requires an independent string of (near-)perfect randomness as an initial resource, which is difficult to justify and leads to circularity: one needs near-perfect randomness to generate more of it. Two-source extraction relaxes this requirement, allowing the second string to be only weakly random, in the sense that it has some min-entropy and/or only a relaxed notion of independence (for example, the case of using a two-source extractor secure in the Markov model Arnon-Friedman et al. (2015), where the two input sources can be correlated through a common cause, or if the sources may have bounded coordination, cross-influence, or mutual information Ball et al. (2022))—although, in this work, we calculate our two-source extractor parameters based on standard independence between the two input sources. Two-source extractors can be used as seeded extractors, simply by assuming that one of the input strings is already near-perfect and independent; therefore, level 3 is strictly a relaxation of the assumptions of level 2.
{Definition}
[…]
$\Delta(\mathsf{Ext_{2}}(X,Y),\mathrm{U}_{m})\leq\epsilon\ ,$ (12)
where $\mathrm{U}_{m}$ is the uniform variable on $\{0,1\}^{m}$. In other words, a two-source extractor is a weakly randomised function that maps a random variable $X$ to a new variable $\mathsf{Ext_{2}}(X,Y)$ that is near-perfect.
{Definition}
[Strong two-source extractor] A two-source randomness extractor is said to be strong in the input $Y$ if the function $\mathsf{Ext_{2}}$ is such that
$\Delta([\mathsf{Ext_{2}}(X,Y),Y],[\mathrm{U}_{m}Y])\leq\epsilon\ ,$ (13)
where [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] denotes the concatenation of random variables and $\mathrm{U}_{m}$ is the uniform variable on $\{0,1\}^{m}$. Strong two-source extractors, like strong seeded extractors, allow for one input source to be used in multiple extraction rounds. The set-up for seeded extraction (implemented using a strong seeded extractor) is shown in Figure 7.
Figure 7: The set-up for two-source extraction. In this case, the initial output of the RNG only needs to have some min-entropy and extraction requires an additional bit string that is weakly random only in the sense that it also has min-entropy.
From the ${\sf{Cryptomite}}$ library, we again use the Circulant extractor Foreman et al. (2024), but this time as a strong two-source extractor. Other extractors in ${\sf{Cryptomite}}$ can be used too, but, since the Circulant extractor offers the best parameters and efficiency, we use it in our implementation. For full details, we refer the reader to Foreman et al. (2024). Two-source extraction requires a second input source with min-entropy above some threshold based on the specific two-source extractor construction. For the Circulant extractor, this requirement is that the sum of the min-entropy rates of the two weak inputs is at least 1. $X_{i}$ is the initial RNG output blocks and $Y$ is the additional weakly random input (which we sometimes call the weak seed) and, as in level 2, we set $|Y|=|X_{i}|+1=$ 10,007.
To generate $Y$, we again use the NIST Randomness Beacon, but, in this case, we minimise the amount of entropy that we assume that it contains, instead of assuming that it has full entropy as in level 2. This change in the assumption increases the likelihood that the assumption holds in practice. The output length of the Circulant extractor is roughly $(\alpha_{\mathsf{NIST}}+\alpha_{\mathsf{RNG}}-1)|Y|$, which we impose by adjusting the min-entropy rate assumption of the NIST Randomness Beacon as $\alpha_{\mathsf{NIST}}$, as
$\alpha_{\mathsf{NIST}}=1.02-\alpha_{\mathsf{RNG}}\,$ (14)
where $\alpha_{\mathsf{RNG}}$ is the min-entropy rate of the initial RNG (found in Section 4). We use 1.02 instead of 1 to account for spurious terms in the parameter calculation that reduce the output length; see Foreman et al. (2024) for the explicit calculation of these penalty terms. In other words, we use the computed min-entropy rate of the RNG under study to minimise the assumption about the second source’s min-entropy rate, whilst imposing a non-trivial output length from the extractor.
The processed output is then generated in two steps: (1) using the Circulant extractor as a two-source extractor on the two input strings $X_{1}$ and $Y$, we generate a (near-)perfect output, which will be the seed in the next step; (2) we use this seed in multiple Circulant seeded extractions on $X_{i\geq 2}$. The multiple outputs of the seeded extractor are concatenated together to obtain a final output of $5\times 10^{10}$ bits. In other words, the concatenation of the two-source and seeded extractors together forms a two-source extractor with advantageous parameters. Therefore, the final output for statistical testing is given by
$\mathsf{Output}$ $=\left[\mathsf{Ext_{s}^{Circulant}}(X_{2},S),\mathsf{Ext_{s}^{Circulant}}(X_{3},S),\ldots,\mathsf{Ext_{s}^{Circulant}}(X_{n},S)\right]\ ,$ (15)
where $S=\mathsf{Ext_{2}^{Circulant}}(X_{1},Y)$, [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] denotes the concatenation of random variables and the extractor round with input $X_{i}$ has error $\epsilon_{\mathsf{ext_{i}}}$. The total error of the final output is $\epsilon_{\mathsf{total}}=\epsilon_{\mathsf{est}}+\epsilon_{\mathsf{ext_{1}}}+\sum_{j=2}^{n}\epsilon_{\mathsf{ext_{j}}}\leq 2^{-32}$. Proof that a strong two-source extractor and strong seeded extractor can be composed into $\mathsf{Ext_{s}^{Circulant}}(X_{i>1},S)$, for $S$ the output of a two-source extractor (right-hand side of Equation 15) can be found in Vadhan (2012) Section 6.3. This, combined with the fact that the composed output error is the sum of each of the individual extractor errors (in Foreman et al. (2024)), allows us to calculate $\epsilon_{\mathsf{total}}$.
@@ 5.2.3 Level 3: Two-Source Extraction
Our results in Table 7 show that all RNGs extracted at level 3 are successful from a statistical perspective, as in the seeded extraction case (level 2). In the Appendices, we implement a variant of level 3 (two-source extraction) where all input strings are drawn from the initial RNG and there is no randomness from an alternative RNG, i.e., rewriting the $\mathsf{Output}$ in Equation 15 using $Y=X_{0}$, where $X_{0}$ is another output block from the initial RNG. In this regime, for near-perfect randomness to be generated, all blocks produced by the initial RNG must be independent of one another (as well as having block min-entropy). Even in this case, the results were successful statistically. Full explanations and results can be found in Section E.1.
Table 7: This table gives the sum of statistical tests failed for $5\times 10$ Gbit samples from each RNG, after strong two-source extraction, taking the RNG as one weak source and randomness from the NIST Randomness Beacon as the second. In cells with multiple entries, failed tests are on the left and suspicious tests (when applicable) are on the right in parentheses. Full results can be found in Section E.2.
{adjustwidth}
[…]
5.2.4 Level 4: Physical Randomness Extraction
@@
Two-source extraction (level 3) allows for the generation of near-perfect randomness if two weakly random but independent strings of randomness are available. In the final level, we consider post-processing with a physical randomness extractor. This level is called physical because it requires a quantum device, in addition to the initial RNG, while the other levels only require mathematical algorithms to perform extraction. At a high level, the role of this additional hardware is to provide a second string of random numbers, whilst making minimal assumptions only.
@@
Adding quantum hardware may initially seem to imply introducing numerous assumptions; however, following the device-independent approach, this hardware can, in principle, be treated as an untrusted black box (which could even have been built by an adversary, so long as it can be shielded once in use and meets some minimal requirements). We call the added assumptions minimal because they are either fundamental to physics—e.g., information cannot travel faster than light speed—or no cryptography can ever be performed without them—e.g., the devices are shielded (there are no backdoors). This is made possible by the development of device-independent protocols, which rely on Bell tests Brunner et al. (2014). The idea is to use the initial RNG to generate random challenges for the quantum device and then study its response. With ideal (noiseless) devices, this approach can be used to self-test the inner functioning of the device, i.e., one can uniquely identify the implemented quantum states and measurements from the observed challenge–response statistics alone. For real (noisy) devices, this approach can be used to bound the adversary’s guessing power, and thus guarantee min-entropy, over the device’s outputs or responses. This approach crucially relies on quantum resources, which have this self-testing property, and has no classical analogue. For a review on the subject, together with its minimal assumptions (called loopholes), we refer the reader to Acín and Masanes (2016). See Figure 8 for an illustration.
@@
Today, quantum devices that are capable of running device-independent protocols are extremely difficult to build (they require the ability to perform a loophole-free Bell Test Brunner et al. (2014)) and exist as experiments on lab benches only. Because of this, more practical implementations have been developed in which a few well-justified assumptions are added (to the minimal ones). The resulting protocols have comparatively fewer assumptions than standard hardware, but not fundamentally minimal. Such a semi-device-independent protocol is the one that we implement for our physical extraction method at level 4, based on an adaptation of the randomness amplification protocol described in Foreman et al. (2023) and implemented on quantum computers. For clarity, the assumptions that we make are the following.
@@
  • The quantum device is assumed to perform a faithful Bell test. This assumption is well motivated when using particular types of devices, such as the quantum computers based on ion traps that we use—see the discussion in Foreman et al. (2023) (Section 6.2 , Validity of quantum computers for Bell experiments and added assumptions).
Figure 8: The set-up for level 4: physical randomness extraction. The initial RNG is used twice: first to generate challenges to the quantum device and second to provide an extra bit string as input to a two-source extractor. The role of the quantum device is to provide an additional source of randomness. The device-independent protocol is performed by using the challenge–response behaviour of the device to obtain a lower bound on the amount of randomness in the device’s responses (without characterising the device itself). The second bit string of the initial RNG and the responses from the quantum device form the two input strings to a two-source extractor, implemented as in level 3.
@@
We used the $\mathsf{H}$1-1 Quantinuum ion-trap quantum computer Quantinuum (1-4 Nov, 2021) as our device to obtain, from its output, a weakly random bit string size of $3.6\times 10^{6}$ bits (this means that, due to using the Circulant extractor, the input length of the initial RNG block to the two-source extraction is also $3.6\times 10^{6}$ bits) with min-entropy rate $\alpha_{\mathsf{Q}}\geq 0.518$, certified in the semi-device-independent manner described above. The Circulant extractor requires $\alpha_{\mathsf{Q}}+\alpha_{\mathsf{RNG}}>1$ to give a non-vanishing output, implying that the rate of an initial RNG must satisfy $\alpha_{\mathsf{RNG}}>0.482$ to allow for physical extraction with our implementation. (This minimum requirement is particularly interesting, since, even if one has access to two identical (and independent) copies of an initial RNG with $\alpha_{\mathsf{RNG}}=0.482+\delta$ for $\delta\in(0,0.18)$, one would be unable to extract from the two (step 3) with today’s implemented extractors. Note that this is not a fundamental limitation, as other two-source extractors allow for one of the strings to have a logarithmic min-entropy rate only. However,  to our knowledge no such extractor has been implemented, let alone efficiently). The advantage of using a quantum device, and therefore level 4, is two-fold: (a) one obtains a rigorous, semi-device-independent lower bound on a second bit string’s min-entropy and (b) the min-entropy rate of the quantum device is above $0.5$, allowing extraction from a weak initial output with rate $0.5$ using the Circulant extractor. Note that the min-entropy of the LSFR was too low to perform physical extraction (its min-entropy rate was below $0.482$; see Table 4).
@@
The processed output is then generated in two steps. (1) We generate a (near-)perfect seed using the Circulant extractor as a two-source extractor on the two input strings $X_{1}$, from the initial RNG, and $Y$, from the $\mathsf{H}$1-1 Quantinuum quantum computer. (2) We use this seed in multiple Circulant seeded extractions on $X_{i\geq 2}$, which are concatenated together to obtain a final output of $5\times 10^{10}$ bits. In other words, the concatenation of the two-source and seeded extractors together again forms a two-source extractor with advantageous parameters. Therefore, the final output for statistical testing is given by
$\mathsf{Output}=\left[\mathsf{Ext_{s}^{Circulant}}(X_{2},S),\mathsf{Ext_{s}^{Circulant}}(X_{3},S),\ldots,\mathsf{Ext_{s}^{Circulant}}(X_{n},S)\right]\ ,$ (16)
@@
where $S=\mathsf{Ext_{2}^{Circulant}}(X_{1},Y)$, [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] denotes concatenation and the extractor round with input $X_{i}$ has error $\epsilon_{\mathsf{ext_{i}}}$. The total error of the final output is $\epsilon_{\mathsf{total}}=\epsilon_{\mathsf{est}}+\epsilon_{\mathsf{ext_{1}}}+\sum_{j=2}^{n}\epsilon_{\mathsf{ext_{j}}}\leq 2^{-32}$. This last step is similar to that of level 3, where the NIST Randomness Beacon is replaced by the $\mathsf{H}$1-1 Quantinuum quantum computer. The statistical test results are given in Table 8.
@@
Table 8: This table gives the sum of statistical tests failed for $5\times 10$ Gbit samples from level 4. Note: The 32-bit LFSR does not generate any output in this setting, since its min-entropy is too low for extraction. In cells with multiple entries, failed tests are on the left and suspicious tests (when applicable) are on the right in parentheses. Full results can be found in Appendix F.
{adjustwidth}
[…]
@@
The results of the statistical tests in Table 8 show, as for levels 2 and 3, that the post-processed RNGs perform well at level 4.
6 Discussion
In this work, we have presented a variety of extraction methods to post-process the output of random number generators (RNGs) and evaluated their impacts on the statistical properties of three widely used RNGs. We started by extensively testing the output from three RNGs and identified statistical failures in two of them, corroborating and extending previous findings Hurley-Smith and Hernandez-Castro (2020); Hamburg et al. (2012). For the RNGs that failed, all post-processing methods improved the statistical properties. Specifically, we found that the processed outputs processed with level 2 or higher (seeded, two-source, or physical extraction) were statistically indistinguishable from uniform distributions. However, due to the inherent limitations of statistical testing, we could not identify any examples that failed under level 2 post-processing (seeded extraction) but succeeded with level 3 or higher (two-source and physical extraction), even though level 3 is provably stronger than level 2.